# Dynamical phase transitions in the non-reciprocal Ising model

### v1

@@ Appendix A The effect of different update rules
For fixed coupling parameters ($\tilde{J}$ and $\tilde{K}$), the different update rules changed the dynamics at steady state. However, they did not lead to qualitative changes in the phase diagram. That way, the critical point $\tilde{J}_{c}$, for some fixed $\tilde{K}$, is altered, but not the existence / destabilization of a phase transition. As an example, we show in Fig. 37 that the phase transition from disorder to swap in 3D, with random sequential updates (standard Glauber algorithm), is qualitatively the same as with the checkerboard algorithnm which was used to produced Fig. 13.
We did not find any qualitative difference related to phase transitions. The only notable difference is that with synchronous updates, fluctuations were considerably larger due to “checkerboard patterns” (small regions in which each spin is anti-aligned with its same-species neighbours) that frequently appeared and then disappeared.
@@
Appendix B Selfish free energy
The dynamics of non-conserved fields $m_{\alpha}(\vec{r},t)$ with $\alpha=1,...,N$ (model A) in equilibrium is [200, 27, 201]
$\partial_{t}m_{\alpha}(\vec{r},t)=-\mu_{\alpha}\frac{\delta F[m_{1},...,m_{N}]}{\delta m_{\alpha}(\vec{r},t)}+\eta_{\alpha}(\vec{r},t)$ (66)
where $F$ is the effective free energy functional, $\mu_{\alpha}$ are generalized mobilities, and $\eta_{\alpha}$ is a white noise satisfying
[…]
where $D_{\alpha}=\mu_{\alpha}k_{B}T$.
For non-reciprocal systems, there is no global $F$ generating the dynamics. However, we extend the notion of a free energy by writing the dynamics of the two magnetization fields $m_{A}$ and $m_{B}$ as
$\partial_{t}m_{\alpha}(\vec{r},t)=-\mu_{\alpha}\frac{\delta F_{\alpha}[m_{A},m_{B}]}{\delta m_{\alpha}(\vec{r},t)}+\eta_{\alpha}(\vec{r},t)$ (68)
where $F_{\alpha}$ is the selfish free energy of the $\alpha$-species. We set $\mu_{\alpha}=1$ and use Eq. (28) to obtain the selfish free energies of the NR Ising model in the fully anti-symmetric case,
$F_{A}$ $=\frac{1}{2}\left(1-\tilde{J}\right)m_{A}^{2}-\tilde{K}m_{A}m_{B}+D\left(\nabla m_{A}\right)^{2}$
$+\frac{1}{12\tilde{J}}\left(\tilde{J}m_{A}+\tilde{K}m_{B}\right)^{4}$ (69)
$F_{B}$ $=\frac{1}{2}\left(1-\tilde{J}\right)m_{B}^{2}+\tilde{K}m_{A}m_{B}+D\left(\nabla m_{B}\right)^{2}$
$+\frac{1}{12\tilde{J}}\left(\tilde{J}m_{B}-\tilde{K}m_{A}\right)^{4}.$
Equation (B) can be viewed as the free-energy landscape that each species “sees”, where the free-energy landscape of $A$ changes with $m_{B}$ and vice versa.
We now restrict ourselves to the zero-dimensional case ($D=0$), for simplicity. In order to see the free-energy landscape of each species, we solve the MF equation (Eq. (28) with $D=0$) and substitute the $m_{B}(t)$ solution into $F_{A}$ and the $m_{A}(t)$ solution into $F_{B}$. The free-energy landscape in steady state is shown in Fig. 38, for the disordered, swap, and static-order phases. The actual values of $m_{A}$ and $m_{B}$ in steady state are shown as disks moving on the free-energy landscape.
Fig. 38: Selfish free-energy landscape. $F_{A}$ as a function of $m_{A}$ (blue) and $F_{B}$ as a function of $m_{B}$ (red) in zero-dimension NR Ising model. $m_{A}(t)$ and $m_{B}(t)$, obtained numerically by solving Eq. (28) with $D=0$ at steady state, are substituted into $F_{A}$ and $F_{B}$ defined in Eq. (B), respectively. Actual $m_{A}$ and $m_{B}$ values are shown as disks. (a) Disorder, $\tilde{J}=0.9$, (b) swap, $\tilde{J}=1.11$, and (c) static-order, $\tilde{J}=1.3$, and in all plots $\tilde{K}=0.1$. In panel a, $A$ and $B$ lines as well as disks are identical and therefore only B is shown.
In the disordered phase (Fig. 38a), both $m_{A}$ and $m_{B}$ are in the global minimum of their respective free energies. In the swap phase (Fig. 38b), the free-energy landscape of each species changes continuously. While one species is momentarily in a global minimum, the other species is descending along the free-energy landscape which changes the free-energy landscape of the first species and so on. In the static order phase (Fig. 38c), one of the species is at its global minimum while the second is in a local minimum, that is not the global one. This configuration is stable in MF, but once noise is introduced, it becomes metastable: the second species will “jump” over the barrier and reach global minimum, making the first species metastable and the cycle repeats itself. In a spatially extended system, this transition happens via droplet nucleation and growth as depicted in Sec. IX.
Appendix C Simulations of the MF equation in 2D
@@
Fig. 39: The effect of added noise on the MF equation in 2D. Simulations of Eq. (70) in 2D comparing the cases of no added noise with added noise and ordered vs. disordered initial conditions. First column is kymographs of the $\theta$ field for a single row on the 2D system. Second, third, and forth columns are snapshots of the $m_{A}$, $m_{B}$, and $\theta$ fields, respectively, taken after the system has reached steady state. Without added noise, the steady state is a homogeneous swap phase for ordered initial conditions and an inhomogeneous spiral phase for disordered initial conditions. With enough added noise, the swap phase is unstable even with ordered initial conditions and spiral defects are formed for both initialization types. Simulation time step is $\Delta t=0.05$, noise amplitude for the case of added noise is $\zeta=0.36$, and other system parameters are: $\tilde{J}=2$, $\tilde{K}=0.8$, $L=200$, and $\tau=1$. At $t=0$, we set $m_{A},m_{B}=1\pm 0.1$ for ordered initial conditions and $m_{A},m_{B}=\pm 0.1$ for disordered initial conditions, where the plus/minus sign is chosen at random for each spin. Kymographs show evolution between $t=9900$ and $t=10000$.
@@
We simulate the full MF equation on a grid (Eq. (21)) in 2D in the regime corresponding to a stable swap phase, and in the fully anti-symmetric case. Results are shown in Fig. 39. Depending on initial conditions, the system can end up in different states. An initially ordered state leads to a stable swap phase where the $\theta$ field changes periodically and homogeneously (Fig. 39, first row), while an initially fully disordered state ($m_{A}$ and $m_{B}$ are chosen randomly at each point to be either $+0.1$ or $-0.1$) reaches a stable state with spiral defects, whose cores barely move over time (Fig. 39, second row).
@@
We test the effect of adding noise to the MF equation by writing,
$\partial_{t}m_{i}^{\alpha}=-\frac{m_{i}^{\alpha}}{\tau}+\frac{1}{\tau}\tanh\left(\frac{J\sum_{j\,{\rm nn\,of}\,i}m_{j}^{\alpha}+K\varepsilon_{\alpha\beta}m_{i}^{\beta}}{k_{B}T}\right)+\eta_{i}^{\alpha}.$ (70)
@@
where $\eta_{i}^{\alpha}$ is white noise satisfying $\langle\eta_{i}^{\alpha}\rangle=0$ and $\langle\eta_{i}^{\alpha}\left(t\right)\eta_{j}^{\beta}\left(t^{\prime}\right)\rangle=\zeta\delta_{\alpha\beta}\delta_{ij}\delta\left(t-t^{\prime}\right)$. Above some critical $\zeta$, and for fixed system size, we observed that even if the system is initialized in an ordered state, after a long time, spiral defects form (see Fig. 39, third row): they are qualitatively similar to those observed when random initial conditions are implemented (Fig. 39, fourth row). In that sense, the MF equation with noise exhibits the same qualitative behavior as the Monte-Carlo simulations, where the swap phase is unstable in 2D.
Appendix D 1D NR Ising model
Fig. 40: 1D NR Ising model. (a) $R$ and (b) $\mathcal{L}$ as a function of $\tilde{J}$ for $\tilde{K}=0.3$ and different system size $L$. (c) Kymographs of $\theta$ for $\tilde{J}=3$ (left) and $\tilde{J}=4$ (right), comparing the Ising model ($\tilde{K}=0$, top) with the NR Ising model ($\tilde{K}=0.2$, middle and $\tilde{K}=0.5$, bottom). System size is $L=500$. Obtained from MC simulations.
The NR Ising model does not have a phase transition to an ordered state (either static or time dependent), as is evident from the decrease of $R$ and $\mathcal{L}$ as system size increases. Similarly to the 1D Ising model, as $L$ increases, domains of opposite magnetization emerge and the system becomes disordered. However, in contrary to the Ising model, the boundaries between the domains do not diffuse but move with non-zero velocity (see Fig. 40c where the kymograph slope indicates the velocity). Domains expand in a way that can be viewed as the 1D version of droplet growth discussed in Sec. IX.
Appendix E 2D numerical phase diagram in the general asymmetric case
@@
Figure 41 shows color maps of $R$ and $\mathcal{L}$ as a function of $\tilde{J}$ and $\tilde{K}_{-}$ with fixed non-zero $\tilde{K}_{+}$, for a 2D system with $L=20$ (panels a-b) and $L=80$ (panels c-d). While for $L=20$ the phase diagram is similar to the MF prediction (Fig. 32), at $L=80$ both $R$ and $\mathcal{L}$ diminish and there is no apparent swap phase, supporting the conclusions of Sec. XII.2.
@@
Fig. 41: 2D numerical phase diagram in the general asymmetric case with fixed (non-zero) $\tilde{K}_{+}$. (a) $R$ and (b) $\mathcal{L}$ color maps as a function of $\tilde{J}$ and $\tilde{K}_{-}$, with $\tilde{K}_{+}=0.1$ and $L=20$. (c-d) The same with $L=80$. Obtained from MC simulations.
Appendix F Additional cuts through the 3D numerical phase diagram in the general asymmetric case
In the general asymmetric case, the phase diagram is three-dimensional as it depends on $\tilde{J}$, $\tilde{K}_{+}$, and $\tilde{K}_{-}$. In the main text the order parameters were calculated as a function of $\tilde{J}$ and $\tilde{K}_{-}$ for a fixed value of $\tilde{K}_{+}$ (Fig. 34). Figure 42 shows additional cuts, both for fixed $\tilde{J}$ and for fixed $\tilde{K}_{-}$, for a system of linear size $L=80$. In both cuts, the three phases of disorder ($R=0$ and $\mathcal{L}=0$), swap ($R\neq 0$ and $\mathcal{L}\neq 0$) and static-order ($R\neq 0$ and $\mathcal{L}=0$) are clearly shown. However, the transition between the swap phase and the static-order phase is mediated by a disordered phase, as was also pointed out in Sec. XII.4. The only region where a direct transition seem to appear is in the $K_{+}\to 0$ limit, but in this case, the static-order phase is supposed to be unstable due to droplets as discussed in Sec. IX, and therefore this direct transition is only observed at finite systems.
@@ Appendix F Additional cuts through the 3D numerical phase diagram in the general asymmetric case
Fig. 42: Additional cuts through the 3D phase diagram in the general asymmetric case. (a) $R$ and (b) $\mathcal{L}$ color maps as a function of $\tilde{K}_{+}$ and $\tilde{K}_{-}$ with $\tilde{J}=1.5$ and $L=80$. (c) $R$ and (d) $\mathcal{L}$ color maps as a function $\tilde{K}_{+}$ and $\tilde{J}$ with $\tilde{K}_{-}=0.3$ and $L=80$. Obtained from MC simulations.
Appendix G Supplementary movies
@@
Movie 2: 3D stable swap phase at steady state. Obtained from MC simulations with parameters: $L=80$, $\tilde{J}=1.5$, and $\tilde{K}=0.1$.
@@
Movie 5: Droplet-capture mechanism in the case of general asymmetric interaction in a 2D system. Obtained from MC simulations with parameters: $L=150$, $\tilde{J}=3.3$, $\tilde{K}_{+}=0.1$, and $\tilde{K}_{+}=0.3$.

### v2

@@ Appendix C The effect of different update rules
For fixed coupling parameters ($\tilde{J}$ and $\tilde{K}$), the different update rules changed the dynamics in steady state. However, they did not lead to qualitative changes in the phase diagram. That way, the critical point $\tilde{J}_{c}$, for some fixed $\tilde{K}$, is altered, but not the existence / destabilization of a phase transition. As an example, we show in Fig. 38 that the phase transition from disorder to swap in 3D, as characterized by the variation of the order parameter $R$ as a function of $\tilde{J}$, is qualitatively the same when using random sequential updates (standard Glauber algorithm), as it is with the checkerboard algorithm, which was used to produced Fig. 12a.
We did not find any qualitative difference related to phase transitions. The only notable difference is that with synchronous updates, fluctuations were considerably larger due to “checkerboard patterns” (small regions in which each spin is anti-aligned with its same-species neighbors) that frequently appeared and then disappeared.
@@
Appendix D Simulations of the mean-field equation in 2D
@@
Fig. 39: The effect of added noise on the mean-field equation in 2D. Simulations of Eq. (91) in 2D comparing the cases of no added noise with added noise and ordered vs. disordered initial conditions. First column is kymographs of the $\theta$ field for a single row on the 2D system. Second, third, and forth columns are snapshots of the $m_{A}$, $m_{B}$, and $\theta$ fields, respectively, taken after the system has reached steady state. Without added noise, the steady state is a homogeneous swap phase for ordered initial conditions and an inhomogeneous spiral phase for disordered initial conditions. With enough added noise, the swap phase is unstable even with ordered initial conditions and spiral defects are formed for both initialization types. Simulations were performed using the Euler–Maruyama method with a time step $\Delta t=0.05$. Noise amplitude for the case of added noise is $\zeta=0.36$, and other system parameters are: $\tilde{J}=2$, $\tilde{K}=0.8$, $L=200$, and $\tau=1$. At $t=0$, we set $m_{A},m_{B}=1\pm 0.1$ for ordered initial conditions and $m_{A},m_{B}=\pm 0.1$ for disordered initial conditions, where the plus/minus sign is chosen at random for each spin. Kymographs show evolution between $t=9900$ and $t=10000$.
@@
We simulate the full mean-field equation on a grid (Eq. (21)) in 2D in the fully anti-symmetric case and in the regime corresponding to a stable swap phase (see Fig. 4). Results are shown in Fig. 39. Depending on initial conditions, the system can end up in different states. An initially ordered state leads to a stable swap phase where the $\theta$ field changes periodically and homogeneously (Fig. 39, first row), while an initially fully disordered state ($m_{A}$ and $m_{B}$ are chosen randomly at each point to be either $+0.1$ or $-0.1$) reaches a stable state with spiral defects, whose cores barely move over time (Fig. 39, second row).
@@
We test the effect of adding noise to the mean-field equation by writing,
$\frac{{\rm d}m_{i}^{\alpha}}{{\rm d}t}=-\frac{m_{i}^{\alpha}}{\tau}+\frac{1}{\tau}\tanh\left(\frac{J\sum_{j\,{\rm nn\,of}\,i}m_{j}^{\alpha}+K\varepsilon_{\alpha\beta}m_{i}^{\beta}}{k_{B}T}\right)+\eta_{i}^{\alpha}.$ (91)
@@
where $\eta_{i}^{\alpha}$ is white noise satisfying $\langle\eta_{i}^{\alpha}\rangle=0$ and $\langle\eta_{i}^{\alpha}\left(t\right)\eta_{j}^{\beta}\left(t^{\prime}\right)\rangle=\zeta\delta_{\alpha\beta}\delta_{ij}\delta\left(t-t^{\prime}\right)$. Above some critical $\zeta$, and for fixed system size, we observed that even when the system is initialized in an ordered state, spiral defects eventually form over time. (see Fig. 39, third row). Moreover, they are qualitatively similar to those observed when random initial conditions are implemented (Fig. 39, fourth row). In that sense, the mean-field equation with noise exhibits the same qualitative behavior as the Monte-Carlo simulations, where the swap phase is unstable in 2D.
Fig. 40: Selfish free-energy landscape. $F_{A}$ as a function of $m_{A}$ (blue) and $F_{B}$ as a function of $m_{B}$ (red) in zero-dimension NR Ising model. Obtained by solving numerically the spatially homogeneous Eq. (30) in steady state, and substituting $m_{A}(t)$ and $m_{B}(t)$ into $F_{A}$ and $F_{B}$ defined in Eq. (95), respectively. Actual $m_{A}$ and $m_{B}$ values are shown as disks. (a) Disorder, $\tilde{J}=0.9$, (b) swap, $\tilde{J}=1.11$, and (c) static order, $\tilde{J}=1.3$; in all plots $\tilde{K}=0.1$. In panel a, $A$ and $B$ lines as well as disks are identical and therefore only B is shown.
Appendix E Selfish free energy
The dynamics of non-conserved fields $m_{\alpha}(\vec{r},t)$ with $\alpha=1,...,N$ (model A) in equilibrium can be written as [223, 27, 247]
$\partial_{t}m_{\alpha}(\vec{r},t)=-\mu\frac{\delta\mathcal{F}[m_{1},...,m_{N}]}{\delta m_{\alpha}(\vec{r},t)}+\eta_{\alpha}(\vec{r},t)$ (92)
where $\mathcal{F}$ is the effective free energy functional, $\mu$ is a generalized mobility, and $\eta_{\alpha}$ is a white noise field satisfying $\langle\eta_{\alpha}(\vec{r},t)\rangle=0$ and $\langle\eta_{\alpha}(\vec{r},t)\eta_{\beta}(\vec{r}\,^{\prime},t^{\prime})\rangle=\zeta\delta_{\alpha\beta}\delta(\vec{r}-\vec{r}\,^{\prime})\delta(t-t^{\prime})$ where $\zeta=2\mu k_{B}T$.
For nonreciprocal systems, there is no global $\mathcal{F}$ generating the dynamics. However, we extend the notion of a free energy by writing the dynamics of the two magnetization fields $m_{A}$ and $m_{B}$ as
$\partial_{t}m_{\alpha}(\vec{r},t)=-\mu\frac{\delta\mathcal{F}_{\alpha}[m_{A},m_{B}]}{\delta m_{\alpha}(\vec{r},t)}+\eta_{\alpha}(\vec{r},t)$ (93)
where $\mathcal{F}_{\alpha}$ is the selfish free energy of the $\alpha$-species. We set $\mu=1$ and use Eq. (30) (interpreted as the mean-field version of Eq. (93)) to obtain the selfish free energies of the nonreciprocal Ising model in the fully anti-symmetric case,
$\mathcal{F}_{A}$ $=\int{\rm d}\vec{r}\bigg{[}\frac{1}{2}(1-\tilde{J})\,m_{A}^{2}-\tilde{K}m_{A}m_{B}+\frac{D}{2}\left|\nabla m_{A}\right|^{2}$
$+\frac{1}{12\tilde{J}}\left(\tilde{J}m_{A}+\tilde{K}m_{B}\right)^{4}\bigg{]}$ (94)
$\mathcal{F}_{B}$ $=\int{\rm d}\vec{r}\bigg{[}\frac{1}{2}(1-\tilde{J})\,m_{B}^{2}+\tilde{K}m_{A}m_{B}+\frac{D}{2}\left|\nabla m_{B}\right|^{2}$
$+\frac{1}{12\tilde{J}}\left(\tilde{J}m_{B}-\tilde{K}m_{A}\right)^{4}\bigg{]}.$
Equation (E) can be viewed as the free-energy landscape that each species “sees”, where the free-energy landscape of $A$ changes with $m_{B}$ and vice versa.
We now restrict ourselves to the zero-dimensional case, where the selfish free energies become the functions
$F_{A}$ $=\frac{1}{2}(1-\tilde{J})\,m_{A}^{2}-\tilde{K}m_{A}m_{B}+\frac{1}{12\tilde{J}}\left(\tilde{J}m_{A}+\tilde{K}m_{B}\right)^{4}$ (95)
$F_{B}$ $=\frac{1}{2}(1-\tilde{J})m_{B}^{2}+\tilde{K}m_{A}m_{B}+\frac{1}{12\tilde{J}}\left(\tilde{J}m_{B}-\tilde{K}m_{A}\right)^{4}.$
In order to visualize the free-energy landscape of each species at a given state, we solve the spatially homogeneous mean-field equation (Eq. (30)) and substitute the $m_{B}(t)$ solution into $F_{A}$ and the $m_{A}(t)$ solution into $F_{B}$. Snapshots of the steady-state free-energy landscape are shown in Fig. 40, for the disordered, swap, and static-order phases. The actual values of $m_{A}$ and $m_{B}$ in steady state are shown as disks moving on the free-energy landscape.
In the disordered phase (Fig. 40a), both $m_{A}$ and $m_{B}$ are in the global minimum of their respective free energies. In the swap phase (Fig. 40b), the free-energy landscape of each species changes continuously. While one species is momentarily in a global minimum, the other species is descending along the free-energy landscape which changes the free-energy landscape of the first species and so on. In the static-order phase (Fig. 40c), one of the species is at its global minimum while the second is in a local minimum, that is not the global one. This configuration is stable in mean-field, but once noise is introduced, it becomes metastable: the second species will “jump” over the barrier and reach global minimum, making the first species metastable and the cycle repeats itself. In a spatially extended system, this transition happens via droplet nucleation and growth as depicted in Sec. IX.
Appendix F Color maps of the order parameters in 2D in the general asymmetric case
@@
Figure 41 shows color maps of $R$ and $\mathcal{L}$ as a function of $\tilde{J}$ and $\tilde{K}_{-}$ with a fixed non-zero value of $\tilde{K}_{+}$, for a 2D system with $L=20$ (panels a-b) and $L=80$ (panels c-d). For $L=20$ the phase diagram resembles the mean-field prediction (Fig. 30), which includes a swap phase. However, for $L=80$ both $R$ and $\mathcal{L}$ diminish and there is no apparent swap phase, supporting the conclusions of Sec. XII.2.
@@
Fig. 41: Color maps of the order parameters in 2D in the general asymmetric case with fixed (non-zero) $\tilde{K}_{+}$. (a) $R$ and (b) $\mathcal{L}$ color maps as a function of $\tilde{J}$ and $\tilde{K}_{-}$, with $\tilde{K}_{+}=0.1$ and $L=20$. (c-d) The same with $L=80$. Obtained from MC simulations.
@@
Fig. 42: Color maps of the order parameters in 3D in the general asymmetric case for a fixed value of $\tilde{J}$ and for a fixed value of $\tilde{K}_{-}$. (a) $R$ and (b) $\mathcal{L}$ color maps as a function of $\tilde{K}_{+}$ and $\tilde{K}_{-}$ with $\tilde{J}=1.5$ and $L=80$. (c) $R$ and (d) $\mathcal{L}$ color maps as a function $\tilde{K}_{+}$ and $\tilde{J}$ with $\tilde{K}_{-}=0.3$ and $L=80$. Obtained from MC simulations.
Appendix G Color maps of the order parameters in 3D in the general asymmetric case
In the general asymmetric case, the phase diagram is three-dimensional as it depends on $\tilde{J}$, $\tilde{K}_{+}$, and $\tilde{K}_{-}$. In the main text the order parameters of the 3D system were calculated from MC simulations as a function of $\tilde{J}$ and $\tilde{K}_{-}$ for a fixed value of $\tilde{K}_{+}$ (Fig. 32). Figure 42 shows the order parameters on additional cuts through the three-dimensional phase diagram, both for a fixed value of $\tilde{J}$ and for a fixed value of $\tilde{K}_{-}$, for a 3D system of linear size $L=80$. In both cuts, the three phases of disorder ($R=0$ and $\mathcal{L}=0$), swap ($R\neq 0$ and $\mathcal{L}\neq 0$) and static-order ($R\neq 0$ and $\mathcal{L}=0$) are visible. However, the transition between the swap phase and the static-order phase is mediated by a disordered phase, as was also pointed out in Sec. XII.4. The only region where a direct transition appears to occur is in the ${\tilde{K}}_{+}\to 0$ limit, see the bottom-left part of Fig. 42a-b and the upper-left part of Fig. 42c-d. However, in this case, the static-order phase is predicted to be unstable due to droplet growth as discussed in Sec. IX, and therefore this direct transition is only observed in finite systems.
Appendix H Supplementary movies
@@
Movie 2: 3D stable swap phase in steady state. Obtained from MC simulations with parameters: $L=80$, $\tilde{J}=1.5$, and $\tilde{K}=0.1$.
@@
Movie 5: Droplet-capture mechanism in a 2D system when inter-species interactions are not fully anti-symmetric. Obtained from MC simulations with parameters: $L=150$, $\tilde{J}=3.3$, $\tilde{K}_{+}=0.1$, and $\tilde{K}_{-}=0.3$.